A person wears a hearing aid that uniformly increases the intensity level of all audible frequencies of sound by 30.0 dB. If the hearing aid picks up a sound with an intensity of 3.5 x 10-9 W/m2 , what is the sound intensity delivered to the eardrum?


1
Expert's answer
2022-05-10T13:21:04-0400
"L=10\lg\frac{I}{I_0}"

"I=I_010^{L/10}\\
=3.5*10^{-9}*10^{30.0/10}=3.5*10^{-6}\:\rm W/m^2"